OpenStudy (anonymous):

The function s(t)=t^3/3-6t^2/2+5t+7 describes the position of a particle in motion. When is the particle moving at a constant speed?

OpenStudy (ivancsc1996):

Well first you must find the first dervivative of position which is velocity, then the second one which is acceleration and then solve t for when acceleration is 0, if you give mw time I'll do the maths.

OpenStudy (anonymous):

just wolfram it

OpenStudy (anonymous):

its a word problem

OpenStudy (anonymous):

of the first and second derivative i mean

OpenStudy (anonymous):

and i have the derivatives but i thought i had to do what ivancsc1996 suggested but for when the particle is resting

OpenStudy (ivancsc1996):

\[\frac{ d }{ dt }s=t ^{2}-6t+5=v\]\[\frac{ d }{ dt }v=2t-6=a \rightarrow a=0 \rightarrow t=3\]The particle is moving at constant speed at t=3s.

OpenStudy (anonymous):

so what would i do for when the particle is resting

OpenStudy (ivancsc1996):

When the particle is resting is when v=0\[v=t ^{2}-6t+5 \rightarrow v=0 \rightarrow t=5,1\]The particle is resting at 5s and 1s

OpenStudy (anonymous):

ok thank you

OpenStudy (ivancsc1996):

Well, that's as much as eleventh grade highschool physics can give. I am sorry if I am wrong.

OpenStudy (anonymous):

this is actually for an AP Calc class but my teachers isnt the best